Which of the following will not represent zero: 
1 + 0

Grade:12

3 Answers

Harshit Singh
3 years ago
Dear Student

1 + 0 = 1
Hence, it does not represent zero

Thanks
Jitendra Kumar
3 years ago
Dear Student,
According to the addition property of 0 states that when 0 is added to any number then we get same number.
So, 1 + 0 = 1
Hence, it does not represent zero
Thanks
 
Shraddha Singh
2 years ago
Dear Student,
According to the addition property of zero , when O is added to any number then we get the same number.
So, 1 + 0 = 1
 
Hence, it does not represent Zero
 
Hope it helps!
Thank you